HUBBLE SPACE TELESCOPE DAILY REPORT #1186 PERIOD COVERED: 0700 EST Daily Status Report as of 203/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5479 (Subarcsecond Structures in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of Mark-1376 and UM625. Two observations were received and displayed. They both were reported to look okay. 1.2 Completed Three Sets of FOS/RD, GHRS/1, WF/PC-2 5642 (Observations of the Impact of the Comet Shoemaker-Levy (1993E) and Jupiter) The Faint Object Spectrograph (Red Side), High Resolution Spectrograph (Side One), and the WF/PC-2 made observations to support the Jupiter/ Shoemaker-Levy event. Thirty observations were received and displayed in the OSS. They all were reported to look okay. 1.3 Complete FGS 5657 (Duplicity Among Hyades Stars: Cycle 4 High) The FGSs (TRANS-Mode) were used to make astrometric observations of star VA162 located in the Hyades cluster in an attempt to fine a close binary. The baseline guide star acquisition for these astrometric observations failed to get fine lock and went to single star coarse track. The Stop flag and the Search Radius Limit Exceeded Flag appeared for FGS#2 during the guide star acquisition. (See Note 2.1). 1.4 Complete Two Sets of FOS/RD 5456 (The BAL Region Covering Factor in QSOs With Weak Narrow Line (O III) Emissions) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of the active galactic nuclei of nearby QSO Q0003+0146 to study weak, narrow O III emissions and strong Fe II emissions in BAL regions. The observations were executed as scheduled, and the data remains on Science Tape Recorder #1. It will be analyzed later. 1.5 Complete FOS/RD 5095 (Spectra at Lambda <3000 Angstroms for QSOs With Z~2 (FOS #3): Cycle 4) The Faint Object Spectrograph (Red Side) was used to make ultraviolet spectral observations of the absorption lines and high-ionization emission lines in QSO PKS0109+17. The binary search failed to located the target and a status buffer error message was generated. The observation was listed as failed. 1.6 Complete WF/PC-2 5369 (HST Medium Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were executed as scheduled, and the data remains on tape recorder #1. It will be analyzed later. 1.7 Complete GHRS/2 5547 (GHRS Side 2 Detractor Calibrations: Cycle 4) The High Resolution Spectrograph (Side 2) was used to make internal flat calibration images to help check the stability fo the detectors. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 5 AT 202/1948Z FGS #1 and #2 failed Fine Lock Acquisition during a ZOE, and FGS #1 went to Course Track. Scheduled Reacquisitions : 3 Successful : 3 2.2 FHST Updates: Scheduled :11 Successful :11 2.3 Significant HSTARs: None 2.4 Operations Notes: The comet Shoemaker-Levy/planet Jupiter collisions concluded early this morning. Today's press briefings will concentrate on the final results. Aboard the HST, all subsystems continued to operate nominally for the entire twenty-four hour period. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-